what is the molar solubility of AgBr in 0.10M FeBr3? Ksp(AgBr)= 5.0*10^-13
a. 3.9*10^-14M
b. 1.7*10^-12M
c. 1.9*10^-11M
d. 2.7*10^-10M
e. 6.2*10^-12M
Answer #1

Given,

Ksp of AgBr = 5.0 x 10-13

Concentration of FeBr3 = 0.10 M and after dissociation concentration of Br- becomes 0.30 M.

Solubility of AgBr will be equals to [Ag+].

Ag+ = Ksp - 5.0 x 10-13 Ag+ - Br-1= 0.3

Ag+ = 16.6 x 10-13 = 1.7 x 10-12M

Option b is correct.